OpenStudy (anonymous):

Evaluate the following integral: 1/([e^-x] +2) dx

OpenStudy (anonymous):

I know I need to multiply by a special of (e^x/e^x) but from there I get shakey

myininaya (myininaya):

multiply top and bottom by e^x then let u=1+2e^x

myininaya (myininaya):

\[\int\limits_{}^{}\frac{e^x}{1+2e^x} dx\]

myininaya (myininaya):

\[u=1+2e^x=>du=2e^x dx =>\frac{1}{2}du=e^x dx\]

OpenStudy (anonymous):

really? i am impressed

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{2} \frac{du}{u}\]

OpenStudy (anonymous):

learn something new every day!

myininaya (myininaya):

\[\frac{1}{2} \ln|u|+C=\frac{1}{2}\ln(1+2e^x) +C\]

OpenStudy (anonymous):

I messed up on my u. Ahhhhh. That's what happened.

OpenStudy (anonymous):

Thanks for the clearly defined steps, I really appreciated it. : )

myininaya (myininaya):

np

myininaya (myininaya):

is this section on hyperbolic trig functions?

myininaya (myininaya):

it seems some cal classes do not cover hyperbolic trig functions

OpenStudy (anonymous):

I just been taught hyperbolic trig this semester. You're right though t semester I took Calc 2, it wasn't taught.

OpenStudy (anonymous):

*last

myininaya (myininaya):

i guess it depends on the instructor. my instructor do not cover it either.

myininaya (myininaya):

but you sort of don't need to cover it i think

myininaya (myininaya):

its just using algebra tricks

OpenStudy (anonymous):

especially since they are usually covered in a 'here's the definition, and here are a couple properties" fashion

OpenStudy (anonymous):

I like hyperbolic trig functions. If you're familiarwith your basic integration techniques then it's not too bad.

OpenStudy (anonymous):

The 1st time I took Calc2 it was presented as "here's the definition" and then bam expected to know how to do them. It sucked.

OpenStudy (anonymous):

myininaya can you help me with this: http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e6432dd0b8b1f45b4aefab4